A 3-ph, 50 Hz overhead transmission line 100 km long delivers 20 MW at 0.9 p.f lagging and at 110 kV. The resistance and reactance per phase per km are (0.2) Ω and (j0.4) Ω respectively, while capacitive admittance per phase per km is (j2.5 * 10-6) siemen. Using nominal T circuit, Find:

1- Constants A, B, C, and D. 2- Sending end voltage.       

3- Sending end current 4- Sending end power factor   

5- Voltage regulation of this T.L.    6- Transmission line efficiency.

Hint: For T equivalent circuit: A = D = (1+ZY/2)   , B= Z(1+ZY/4)   , C= Y
Answer #1

Gliven- felcoHz, lookm, P= 20MW, V = 110ku, pf - oglag. - Ro Total Resistance R = 0.2 x 100km = 20 XL=04x100 soon. Ys 25x16bCharging current - Ic - JYT, -42.5x604 (655+2 +31593) Il = 0.4+1641 Sending End current Iss Ip+Tc = (10.5-5 504) + (-0.4+]